OpenStudy (anonymous):

Two teams have played each other 14 times. Team A won 9 games and team B won 5 games. Bob offers to bet $6 on team A winning while you bet $4 on team B winning. The winner gets $10. How much money should you expect to lose on such a bet? (Round to the nearest cent.)

OpenStudy (dumbcow):

E(x) = (5/14)*6 - (9/14)*4 = -3/7 So you would expect to lose approx 43 cents